-5(1-5x)+5(-8x-2)=-4x-8x

Answer by sheldonbbtrocks(115)

-5(1-5x)+5(-8x-2)=-4x-8x
1. Distribute the first two terms:
-5+25x-40x-10=-4-8x
2. Collect like terms:
-15x-15=-12x
3. Add 15x to both sides:
-15=3x
4. Divide both sides by 3
-5=x

The answer is [x=-5]